OpenStudy (anonymous):

Could you explain how you get cosb = cos^2b -1

OpenStudy (amistre64):

using ( ) help

OpenStudy (amistre64):

cos(b) = cos^2(b) - 1 or cos(b) = cos^2(b-1)

OpenStudy (anonymous):

cos(b) = cos^2(b) - 1

OpenStudy (amistre64):

we could test it out first to see if that is true: cos(pi) = -1 cos^2(pi) - 1 = 0 i dont see how they are equal unless you need to find the points in which they intersect

OpenStudy (anonymous):

first of all it is not true. take for example \[b=\frac{\pi}{4}\] then what amistre said

OpenStudy (amistre64):

for which "b" values are they equal

OpenStudy (anonymous):

perhaps "solve for b"?

OpenStudy (amistre64):

then its a quadratic

OpenStudy (amistre64):

c^2 -c - 1 = 0

OpenStudy (amistre64):

whrere c = cos(b)

OpenStudy (anonymous):

i think he wants 2 prove:cos(2b)=2cos^2(b)-1

OpenStudy (anonymous):

yeah have fun solving that one. get \[c=\frac{1\pm\sqrt{5}}{2}\] which is no big deal. oh. never mind

OpenStudy (amistre64):

i think he wants a strawberry milkshake, but that doesnt help the problem :)

OpenStudy (anonymous):

well I got an answer before that cos(b) = 2cos^2 (b/2) -1 but I couldnt prove it myself , I got an problem how to get from : sin(l) (1 + cos(b) ) to 4sin(l/2)cos(l/2)cos^2(b/2)

OpenStudy (amistre64):

it not an identity ...

OpenStudy (amistre64):

if it were an identity, the all values for b would equate, and they dont

OpenStudy (anonymous):

I know , I just saw it in a book where he got from first sin(l) (1 + cos(b) ) to 4sin(l/2)cos(l/2)cos^2(b/2) but couldnt find how

OpenStudy (anonymous):

@unniverzal you cannot prove it because it is not true for all values of b. like saying "prove \[2x+1=x-4\] " which is not true for all x. you can "solve for x" but you cannot prove that they are always equal.

OpenStudy (amistre64):

i got no idea where: sin(l) (1 + cos(b) ) to 4sin(l/2)cos(l/2)cos^2(b/2) comes from .....

OpenStudy (amistre64):

good luck with it tho :)

jhonyy9 (jhonyy9):

begin with

jhonyy9 (jhonyy9):

\[\sin ^{2}x +\cos ^{2}x =1\]

jhonyy9 (jhonyy9):

consider b=x